OpenStudy (anonymous):

Find the coordinates of the other endpoint of the segment, given its midpoint and one endpoint. midpoint (2, 6), endpoint (-1, 1)

OpenStudy (anonymous):

\[\frac{-1+x}{2}=2\] x = 5 \[\frac{1+y}{2}=6\] y =11 so other endpoint is (5,11)